OpenStudy (anonymous):

2x^2 = -3x-5 using the quadtratic formula

OpenStudy (anonymous):

First, rewrite the equation as a quadratic on one side of the = . i.e. put it in the form \[ax^2+bx+c=0\]

OpenStudy (anonymous):

where a, b, and c are the constant coefficients.

OpenStudy (anonymous):

Once you have rewritten it in the above form, we can apply the quadratic formula, which Im sure you have in front of you, but is \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]